OpenStudy (anonymous):

(w-2)/(w-8)-(w+1)/(w+8)+(w-104)/(w^2-64)=

myininaya (myininaya):

go polpak!

myininaya (myininaya):

\[\frac{w-2}{w-8}\frac{x+8}{x+8}-\frac{w+1}{w+8}\frac{w-8}{w-8}+\frac{w-104}{(w-8)(w+8)}\]

myininaya (myininaya):

gosh stupid x

myininaya (myininaya):

x=w k?

myininaya (myininaya):

\[\frac{(w-2)(w+8)-(w+1)(w-8)+(w-104)}{(w-8)(w+8)}\]

OpenStudy (anonymous):

\[{(w-2) \over (w-8)} - {(w+1)\over (w+8)} + {(w-104) \over (w^2 - 64)} \]\[={(w-2) \over (w-8)}\cdot {w+8 \over w+8} - {(w+1)\over (w+8)} \cdot {w-8 \over w-8} + {(w-104) \over (w^2 - 8^2)} \]\[={(w-2)(w+8) \over (w^2 - 8^2)} - {(w+1)(w-8)\over (w^2 - 8^2)} + {(w-104) \over (w^2 - 8^2)} \]\[= {(w-2)(w+8) - (w+1)(w-8) + (w-104) \over (w^2 - 8^2)} \]

OpenStudy (anonymous):

etc.

myininaya (myininaya):

multiply those binomials and then combine like terms wand woolah you are done well something might cancel

myininaya (myininaya):

you're so sweet polpak

OpenStudy (anonymous):

Did you see my puzzler?

myininaya (myininaya):

no

myininaya (myininaya):

whats its 20?

OpenStudy (anonymous):

Thank you Polpak